### v2

# Description length of canonical and microcanonical models

Abstract
Non-equivalence between the canonical and the microcanonical ensemble has been shown to arise for models defined by an extensive (i.e. ‘scaling with the size of the system’) number of constraints (e.g. the Configuration Model). Here, we compare the description length of binary canonical and microcanonical models in light of ensemble non-equivalence. Specifically, we consider the description length induced by the Normalized Maximum Likelihood (NML), which consists of two terms, i.e. a model log-likelihood and its complexity. While the effects of ensemble non-equivalence on the log-likelihood term are well understood, its effects on the complexity term have not yet been systematically studied. Here, we find that i) microcanonical models are always more complex than their canonical counterparts, ii) the best-scoring model in terms of description length highly depends on the numerical values of the constraints and iii) the difference between the canonical and the microcanonical description length is strongly influenced by the degree of non-equivalence, a result suggesting that non-equivalence should be taken into account when selecting models. Finally, we compare the NML-based approach to model selection with the Bayesian one in light of our results, showing that the Bayesian description length becomes much more sensitive to the choice of the prior when an extensive number of constraints is involved.
I Introduction
Entropy maximization provides a principled approach to inference, resulting in models that are maximally random given certain constraints. These models have found applications across a wide range of scientific disciplines, such as network science squartini_springer ; cimini19 ; squartini18 and time-series analysis marcaccioli2020maximum ; marcaccioli2020correspondence .
Two distinct formulations of maximum-entropy models are possible depending on how constraints are enforced. Canonical models are defined by requiring the value of the constraints to be satisfied on average. Microcanonical models are defined by requiring the value of the constraints to be satisfied exactly. These two formulations do not necessarily lead to the same description of a given system, i.e. they are, in general, non-equivalent.
Non-equivalence between canonical and microcanonical models can be formally defined in several (under mild conditions, equivalent) ways touchette15 . A particularly convenient definition is the measure-level one, which is rooted in information theory and involves the Kullback-Leibler (KL) divergence between the microcanonical and canonical probability distributions. According to this definition, (non-)equivalence is characterized by a (non-)vanishing relative entropy density as the system size approaches infinity. Interestingly, since the KL divergence coincides with the difference between the canonical and the microcanonical log-likelihood functions, measure-level (non-)equivalence can be inspected by simply considering the asymptotic behaviour of such difference squartini15 . Non-equivalence has been recently shown to arise in models that involve an extensive number of constraints (e.g. the Configuration Model for networks squartini15 ; squartini17 ; qi ) and may have practical implications for the expected macroscopic properties of the system under study (see bruno20 for an example concerning bipartite networks).
Whenever the two approaches are equivalent, they identify the same set of typical configurations. In these cases, preferring one formulation over the other becomes a matter of convenience (e.g. computational efficiency). Whenever the two approaches are non-equivalent, the choice depends on the specific characteristics of the system under study and the nature of the constraints involved. For example, in the context of choosing a null model to be compared with real data for e.g. pattern detection, it has been argued that, when non-equivalence holds, the ensemble choice should be based on a theoretical guiding principle garlaschelli17 ; qi ; squartini15bis ; garlaschelli18 ; garlaschelli16 . In this case, if the data are expected to be noisy, meaning that the measured values of the constraints generally do not match the true (unobservable) ones, one should prefer the canonical version of the null model because it allows for fluctuations in the constraint quantities, while the microcanonical one would paradoxically assign zero probability to the true configuration of the system. Conversely, if, by hypothesis, the measured values of the constraints are not affected by error or noise, the microcanonical model should be preferred.
When there is no prior expectation available about the presence or absence of noise in the empirical values of the constraints, the decision between hard and soft constraints should be based on posterior evidence, i.e. which of the two null models best describes the data. This immediately leads to a problem of model selection between canonical and microcanonical ensembles.
Information theory provides a powerful framework for model selection by offering criteria that capture the trade-off between a model ‘goodness of fit,’ quantified by its log-likelihood, and a model ‘complexity.’ The two historically most popular model selection criteria are the Akaike information criterion (AIC) aic and the Bayesian information criterion (BIC) bic , according to which one should pick the model minimizing
$\text{AIC}=-\ln\mathcal{L}+k$ (1)
and
$\text{BIC}=-\ln\mathcal{L}+\frac{k}{2}\ln V,$ (2)
respectively, where $\mathcal{L}$ is the model likelihood, $k$ is the number of its parameters and $V$ is the sample size. Canonical and microcanonical models defined by the same set of constraints share the same number of parameters. Therefore, according to both AIC and BIC, their complexity is the same, and a comparison based on these criteria would reduce to a mere comparison between their log-likelihood functions. This would naively lead to the conclusion that the microcanonical model is always the best-scoring one, since it has the highest likelihood. However, neither AIC nor BIC can be computed for microcanonical models, since they are derived under regularity assumptions that are not even remotely met by the latter. Moreover, both criteria are asymptotic results derived under the assumption of a finite number of parameters as the system size approaches infinity burnham . This assumption is not met by models showing non-equivalence, whose number of parameters scales with the size of the system.
To overcome these issues, here we consider the Minimum Description Length (MDL) principle grunwald_book ; grunwald2019 ; yamanishi2023 , a family of information theoretical approaches to model selection seeking the model that provides the shortest description of data. Specifically, we focus on the approach based upon the Normalized Maximum Likelihood (NML), or Shtarkov, distribution rissanen96 , our aim being that of investigating the interplay between the difference of log-likelihood functions and that of complexities in light of ensemble non-equivalence. Indeed, while the impact of non-equivalence on the difference between log-likelihoods is quite well understood, its effects on the difference between complexities remain an unexplored topic within this framework. To encompass a broad spectrum of applications, our study focuses on maximum-entropy models suitable for studying systems represented by binary rectangular matrices (e.g., bipartite networks, time series). More precisely, we aim at determining i) whether the influence of non-equivalence persists even when the complexity term is accounted for and ii) if non-equivalent canonical and microcanonical models should, in fact, be considered as different from a model selection perspective. We conclude that the answer to both questions is yes, raising the question of whether non-equivalence can be defined on a model selection level. Moreover, as a consequence of our results, we show that, in a Bayesian context, different priors can result in vastly different description lengths when an extensive number of constraints is enforced.
The rest of the paper is organized as follows. Section II introduces MDL for discrete data and the Normalized Maximum Likelihood (NML) approach to define description lengths. In Section III, we provide a general expression for the difference between the description length of canonical and microcanonical models as a function of the Kullback-Leibler divergence between the two distributions. In Section IV, we explicitly derive and compare the description length of both the canonical and the microcanonical formulations of models defined by one global constraint (i.e. the sum of the entries of a matrix) as well as one-sided local constraints (i.e. the row-specific sums of a matrix). While the first constraint leads to ensemble equivalence, the second set of constraints leads to ensemble non-equivalence. Finally, Section V reviews the relationship between the NML-based and Bayesian approaches when non-equivalence holds.
II The Minimum Description Length principle
The MDL principle embodies the idea that the best model to describe a given dataset is the one providing the shortest description. We now introduce the basic formalism by focusing on the discrete data $\bm{x}\in\mathcal{X}$, where $\mathcal{X}$ represents the space of all possible samples of fixed size $V$. A parametric statistical model $\mathcal{M}$ consists of a family of probability distributions sharing the same functional form, i.e.
=$\mathcal{M}=\{P_{\mathcal{M}}(\bm{x};\bm{\theta})\}_{\bm{\theta}\in\Theta}$ (3)
where $\bm{\theta}$ is the vector of parameters whose values lie in the set $\Theta$.
In order to apply the MDL principle, one needs to compute the description length of the data $\bm{x}$ provided by a model $\mathcal{M}$. By Kraft’s inequality, the optimal number of natural digits (nats) needed to describe the outcome $\bm{x}$ of a probability distribution $P(\bm{x})$ is given by $-\ln P(\bm{x})$. Since, however, $\mathcal{M}$ encompasses many distributions corresponding to different parameter values, a natural choice is that of considering the description length provided by the distribution minimizing the quantity $-\ln P_{\mathcal{M}}(\bm{x};\bm{\theta})$, reading
$\mathcal{L}_{\mathcal{M}}(\bm{x})\equiv P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x})),$ (4)
with $\hat{\bm{\theta}}(\bm{x})$ being the maximum-likelihood (ML) estimator of $\bm{\theta}$, i.e.
$\hat{\bm{\theta}}(\bm{x})=\arg\max_{\bm{\theta}\in\Theta}P_{\mathcal{M}}(\bm{x};\bm{\theta}).$ (5)
However, the distribution $\mathcal{L}_{\mathcal{M}}(\bm{x})$ is affected by (at least) two problems having the same origin: the parameters appearing in its definition are evaluated for each individual data in $\mathcal{X}$. As a consequence, i) $\mathcal{L}_{\mathcal{M}}(\bm{x})$ cannot be a probability distribution with support in $\mathcal{X}$ as it is not properly normalized, and ii) adopting $\mathcal{L}_{\mathcal{M}}(\bm{x})$ may lead to severe overfitting as it is defined a posteriori with respect to the observations.
To solve these problems, one can build a so-called universal distribution relative to $\mathcal{M}$, i.e. a unique, representative distribution $\bar{P}_{\mathcal{M}}(\bm{x})$ of the statistical model $\mathcal{M}$. Once equipped with a universal distribution, the description length of $\bm{x}$ through is $\mathcal{M}$ is defined as
$\text{DL}^{\bar{P}}_{\mathcal{M}}(\bm{x})\equiv-\ln\bar{P}_{\mathcal{M}}(\bm{x}).$ (6)
Universal distributions (which don’t necessarily belong to $\mathcal{M}$) can be defined in several ways. In what follows, we will consider the MDL recipe based upon the NML universal distribution (for a more formal introduction to universal coding, we redirect the interested reader to grunwald_book ).
II.1 The Normalized Maximum Likelihood universal distribution
The NML distribution corresponds to the distribution attaining the minimum point-wise regret in the worst-case scenario grunwald_book , i.e.
$\bar{P}^{\textrm{NML}}_{\mathcal{M}}\equiv\arg\min_{\bar{P}}\max_{\bm{x}\in\mathcal{X}}\left\{-\ln\bar{P}(\bm{x})-[-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))]\right\}.$ (7)
In words, it is the distribution that a priori requires the minimum number of nats when compared to the shortest description length a posteriori, that is $-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))$. The minimum is found in the worst-case scenario, i.e. for the data $\bm{x}$ for which this difference is maximal. The unique solution to this minimax problem reads
$\bar{P}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})=\frac{P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))}{\sum_{\bm{y}\in\mathcal{X}}P_{\mathcal{M}}(\bm{y};\hat{\bm{\theta}}(\bm{y}))},$ (8)
which is a normalized maximum likelihood. Following the universal coding approach, the description length of $\bm{x}$ through model $\mathcal{M}$ is computed as
$\text{DL}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})\equiv$ $-\ln\bar{P}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})$
$=$ $-\ln P_{\mathcal{M}}(\bm{x};\hat{\bm{\theta}}(\bm{x}))+\ln\sum_{\bm{y}\in\mathcal{X}}P_{\mathcal{M}}(\bm{y};\hat{\bm{\theta}}(\bm{y})).$ (9)
While the first term is (minus) a model maximum log-likelihood, the second term
$\text{COMP}_{\mathcal{M}}\equiv\ln\sum_{\bm{y}\in\mathcal{X}}P_{\mathcal{M}}(\bm{y};\hat{\bm{\theta}}(\bm{y}))=\ln\sum_{\bm{y}\in\mathcal{X}}\mathcal{L}_{\mathcal{M}}(\bm{y})$ (10)
is named parametric complexity of $\mathcal{M}$. Notice that $\mathcal{X}$ represents the set of all possible data sets of fixed size $V$, i.e. all the observable samples of size $V$. Thus, $\text{COMP}_{\mathcal{M}}$ depends on both the model and $\mathcal{X}$. Finally, we can rewrite equation (II.1) as
$\text{DL}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})=-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\text{COMP}_{\mathcal{M}}.$ (11)
Similarly to other criteria rooted in information theory, this description length consists of two terms: the log-likelihood term, evaluated on the data and favouring models with a high goodness-of-fit, and the complexity term, depending only on the chosen model and penalizing the ones that are too complex.
Once we have evaluated the description length of our data according to a basket of competing models, the MDL principle prescribes selecting the one providing the minimum DL.
III Description length of canonical and microcanonical models
Although computing the complexity term within the NML-based description length can be challenging, models defined by a sufficient statistic admit a simpler expression for this term. This section introduces an important class of models of this kind, the maximum-entropy models (MEMs), both in their canonical and microcanonical formulations. Specifically, we will focus on discrete data that can be represented as a binary $n\times m$ matrix $\mathbf{G}$, i.e.
$\mathbf{G}=\{g_{ij}\}_{\begin{subarray}{c}i=\{1\dots n\}\\ j=\{1\dots m\}\end{subarray}}$ (12)
and denote by $\mathcal{G}$ the set of all such matrices that, when endowed with a probability distribution $P(\mathbf{G})$, constitutes a proper ensemble of matrices. The matrix representation is often used to describe systems composed of $n$ elements to be modelled (corresponding to the number of rows), each characterized by $m$ state variables or degrees of freedom (corresponding to the number of columns). The size of the data set is fixed and supposed to represent the number of independent entries of the matrix, i.e. $nm$.
In the most common scenario, we can access a vector of quantities $\mathbf{c}^{*}=\mathbf{c}(\mathbf{G}^{*})$, evaluated on the only available matrix $\mathbf{G}^{*}$, and representing the sufficient statistic of the model. Maximum-entropy models are, then, obtained by looking for the probability distribution maximizing Shannon entropy
$\mathcal{S}[P]=-\sum_{\mathbf{G}\in\mathcal{G}}P(\mathbf{G})\ln P(\mathbf{G})$ (13)
while constraining the sufficient statistics. If the model is required to reproduce the value of the constraints exactly, i.e. $\mathbf{c}(\mathbf{G})=\mathbf{c}^{*}$, one obtains a microcanonical model. If, instead, the model is required to reproduce the value of the constraints on average, i.e. $\langle\mathbf{c}(\mathbf{G})\rangle_{P}=\mathbf{c}^{*}$ with $\langle\cdot\rangle_{P}$ being the ensemble average induced by distribution $P$, one obtains the corresponding canonical formulation. This approach, introduced by Gibbs gibbs and reprised by Jaynes jaynes , leads to ensembles of matrices that reproduce (either exactly or on average) the constrained quantities while randomizing everything else.
In what follows, we will adopt the NML-based approach to compute the description lengths of microcanonical and canonical MEMs, which is optimal (in a minimax sense) when assuming no prior knowledge on the parameters. Thus, the description length is taken as the one defined in equation (11). For models admitting a sufficient statistic, the complexity term takes the convenient form
$\text{COMP}=\ln\sum_{\mathbf{c}\in\mathcal{C}}\Omega({\mathbf{c}})\mathcal{L}(\mathbf{c}),$ (14)
where $\mathcal{L}(\mathbf{c})$ denotes the value of the maximum likelihood in correspondence of a sufficient statistics reading $\mathbf{c}$ and $\Omega(\mathbf{c})$ denotes the number of configurations with $\mathbf{c}(\mathbf{G})=\mathbf{c}$, i.e.
$\Omega(\mathbf{c})\equiv\sum_{\mathbf{G}\in\mathcal{G}}\delta_{\mathbf{c}(\mathbf{G}),\mathbf{c}}$ (15)
with $\delta_{i,j}$ being the Kronecker delta. As this set may be empty for some values of $\mathbf{c}$, we will solely focus on the graphical values, i.e. the set $\mathcal{C}=\{\mathbf{c}:\Omega(\mathbf{c})\neq\emptyset\}$ of values that are verified by at least one configuration, and denote it with $\mathcal{N}_{\mathcal{C}}=|\mathcal{C}|$ its cardinality. In other words, the sum in equation (14) runs over the graphical values of the sufficient statistics rather than over all possible data.
III.1 Description length of microcanonical models
When considering microcanonical models, entropy maximization yields the following functional form
$P_{\textrm{mic}}(\mathbf{G};\mathbf{c})=\begin{dcases}\frac{1}{\Omega(\mathbf{c})}&\text{if}\>\mathbf{c}(\mathbf{G})=\mathbf{c}\\ 0&\text{else},\end{dcases}$ (16)
where $\mathbf{c}$ is the vector of parameters whose values lie in the discrete set $\mathcal{C}$ for microcanonical models, in fact, the vector of parameters corresponds to the sufficient statistics itself - and $P_{\textrm{mic}}(\mathbf{G};\mathbf{c})$ is a uniform probability whose mass differs from zero only in correspondence of those configurations that verify the condition $\mathbf{c}(\mathbf{G})=\mathbf{c}$.
We will now evaluate the description length of a microcanonical model. Its maximum log-likelihood simply reads
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{c})=\ln\frac{1}{\Omega(\mathbf{c})}=-\ln\Omega(\mathbf{c})$ (17)
while combining this equation with equation (14), one obtains the following expression for the related complexity:
$\text{COMP}_{\textrm{mic}}=\ln\sum_{\mathbf{c}\in\mathcal{C}}\Omega(\mathbf{c})\cdot\frac{1}{\Omega(\mathbf{c})}=\ln\sum_{\mathbf{c}\in\mathcal{C}}1=\ln\mathcal{N}_{\mathcal{C}}.$ (18)
Hence, the description length of a microcanonical maximum-entropy model is
$\text{DL}_{\textrm{mic}}^{\text{NML}}(\mathbf{G}^{*})=\ln\Omega(\mathbf{c}^{*})+\ln\mathcal{N}_{\mathcal{C}}.$ (19)
III.2 Description length of canonical models
When considering canonical models, entropy maximization yields the following well-known exponential form
$P_{\textrm{can}}(\mathbf{G};\bm{\theta})=\frac{e^{-\bm{\theta}\cdot\mathbf{c}(\mathbf{G})}}{Z(\bm{\theta})}$ (20)
with $Z(\bm{\theta})=\sum_{\mathbf{G}\in\mathcal{G}}e^{-\bm{\theta}\cdot\mathbf{c}(\mathbf{G})}$ being the partition function. If the system under study is a network, the maximum-entropy canonical models are also known as Exponential Random Graph Models (ERGMs). In this case, the vector of parameters $\bm{\theta}$ corresponds to the vector of Lagrange multipliers needed to carry out a constrained entropy maximization. The ML estimators $\bm{\hat{\theta}}(\mathbf{G})$ are found by solving the system of equations
$\langle\mathbf{c}\rangle_{\bm{\hat{\theta}}(\mathbf{G})}=\mathbf{c}(\mathbf{G})$ (21)
with $\langle\cdot\rangle_{\bm{\theta}}$ representing the ensemble average induced by $P_{\textrm{can}}(\mathbf{G};\bm{\theta})$.
The maximum log-likelihood of a canonical model is linked to the microcanonical one through the Kullback-Leibler divergence:
$\left.S(P_{\textrm{mic}}||P_{\textrm{can}})\right|_{\mathbf{c},\bm{\theta}}=\sum_{\mathbf{G}\in\mathcal{G}}P_{\textrm{mic}}(\mathbf{G};\mathbf{c})\ln\frac{P_{\textrm{mic}}(\mathbf{G};\mathbf{c})}{P_{\textrm{can}}(\mathbf{G};\bm{\theta})}.$ (22)
In fact, it can be shown squartini17 ; squartini15 that
$D_{\text{KL}}(\mathbf{c^{*}})$ $\equiv S(P_{\textrm{mic}}||P_{\textrm{can}})|_{\mathbf{c}^{*},\bm{\theta}^{*}}$
$=\ln\frac{P_{\textrm{mic}}(\mathbf{G}^{*};\mathbf{c}^{*})}{P_{\textrm{can}}(\mathbf{G}^{*};\bm{\theta}^{*})}$
$=\ln\mathcal{L}_{\textrm{mic}}(\mathbf{c}^{*})-\ln\mathcal{L}_{\textrm{can}}(\mathbf{c}^{*})$ (23)
where $\bm{\theta}^{*}=\hat{\bm{\theta}}(\mathbf{c}^{*})$. Hence,
$\mathcal{L}_{\textrm{can}}(\mathbf{c})$ $=\mathcal{L}_{\textrm{mic}}(\mathbf{c})\cdot e^{-D_{\text{KL}}(\mathbf{c})}=\frac{1}{\Omega(\mathbf{c})}\cdot e^{-D_{\text{KL}}(\mathbf{c})}$ (24)
for any value of the sufficient statistic, and
$\text{COMP}_{\textrm{can}}^{\text{NML}}$ $=\ln\sum_{\mathbf{c}\in\mathcal{C}}\Omega(\mathbf{c})\cdot\frac{1}{\Omega(\mathbf{c})}\cdot e^{-D_{\text{KL}}(\mathbf{c})}=\ln\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}.$ (25)
As a consequence, the description length of a canonical model can be expressed as
$\text{DL}_{\textrm{can}}^{\text{NML}}(\mathbf{G}^{*})$ $=-\ln\mathcal{L}_{\textrm{can}}(\mathbf{c}^{*})+\ln\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}$
$=\ln\Omega(\mathbf{c}^{*})+D_{\text{KL}}(\mathbf{c}^{*})+\ln\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}.$ (26)
III.3 Comparing canonical and microcanonical models
We can now compare the description length of canonical and microcanonical models. Given that $D_{\text{KL}}$ is guaranteed to be non-negative, we have
$\Delta\ln\mathcal{L}(\mathbf{G}^{*})$ $\equiv\ln\mathcal{L}_{\textrm{can}}(\mathbf{G}^{*})-\ln\mathcal{L}_{\textrm{mic}}(\mathbf{G}^{*})$
$=-D_{\text{KL}}(\mathbf{c}^{*})\leq 0$ (27)
i.e. the canonical likelihood is always smaller than or equal to the microcanonical one. For the same reason, it is straightforward to see that the canonical complexity is always smaller than, or equal to, the microcanonical one:
$\Delta\text{COMP}$ $\equiv\text{COMP}_{\textrm{can}}-\text{COMP}_{\textrm{mic}}$
$=\ln\frac{\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{KL}(\mathbf{c})}}{\mathcal{N}_{\mathcal{C}}}\leq 0.$ (28)
Thus, microcanonical models achieve a higher goodness-of-fit but, at the same time, are more complex. The interplay between the two differences determines the difference between the canonical and the microcanonical description length:
$\Delta\text{DL}(\mathbf{G}^{*})$ $\equiv\text{DL}^{\text{NML}}_{\textrm{can}}(\mathbf{G}^{*})-\text{DL}^{\text{NML}}_{\textrm{mic}}(\mathbf{G}^{*})$
$=-\Delta\ln\mathcal{L}(\mathbf{G}^{*})+\Delta\text{COMP}$
$=D_{\text{KL}}(\mathbf{c}^{*})+\ln\frac{\sum_{\mathbf{c}\in\mathcal{C}}e^{-D_{\text{KL}}(\mathbf{c})}}{\mathcal{N}_{\mathcal{C}}}.$ (29)
The following reasoning provides a different way of looking at the expression above. The canonical distribution $P_{\text{can}}(\mathbf{G};\bm{\theta})$ induces the following distribution on the values of the sufficient statistic
$Q_{\text{can}}(\mathbf{c};\bm{\theta})=\Omega(\mathbf{c})P_{\text{can}}(\mathbf{c};\bm{\theta})$ (30)
where $P_{\text{can}}(\mathbf{c};\bm{\theta})$ is the canonical probability of a matrix $\mathbf{G}$, such that $\mathbf{c}(\mathbf{G})=\mathbf{c}$, for a given value of the vector of parameters $\bm{\theta}$. The maximum likelihood assigned to $\mathbf{c}$ by the correspondent model $\{Q_{\text{can}}(\mathbf{c},\bm{\theta})\}_{\theta\in\Theta}$ is, thus,
$Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))$ $=\Omega(\mathbf{c})P_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\bm{c}))=\frac{\mathcal{L}_{\text{can}}(\mathbf{c})}{\mathcal{L}_{\text{mic}}(\mathbf{c})}=e^{-D_{\text{KL}}(\mathbf{c})}$ (31)
and
$\Delta\text{DL}(\mathbf{G}^{*})=-\ln Q_{\text{can}}(\mathbf{c}^{*};\bm{\theta}^{*})+\ln\frac{\sum_{\mathbf{c}\in\mathcal{C}}Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))}{\mathcal{N}_{\mathcal{C}}};$ (32)
upon defining
$\bar{Q}_{\text{can}}\equiv\frac{\sum_{\mathbf{c}\in\mathcal{C}}Q_{\text{can}}(\mathbf{c},\hat{\bm{\theta}}(\mathbf{c}))}{\mathcal{N}_{\mathcal{C}}}$ (33)
as the uniform average of $Q_{\text{can}}(\mathbf{c};\hat{\bm{\theta}}(\mathbf{c}))$ over all possible values of $\mathbf{c}\in\mathcal{C}$, we find that the difference between description lengths can be expressed as
$\Delta\text{DL}(\mathbf{G}^{*})=\ln\frac{\bar{Q}_{\text{can}}}{Q_{\text{can}}^{*}}$ (34)
where $Q_{\text{can}}^{*}\equiv Q_{\text{can}}(\mathbf{c}^{*},\bm{\theta}^{*})$. This expression depends only on the canonical model, providing a new interpretation of the comparison between the canonical and the microcanonical description length: the canonical model provides a shorter description length whenever $\Delta\text{DL}(\mathbf{G}^{*})<0$, i.e. whenever $Q_{\text{can}}^{*}>\bar{Q}_{\text{can}}$. In other words, the canonical model is to be preferred over the microcanonical one whenever the maximum likelihood it assigns to the observed value $\mathbf{c}^{*}$ is higher than the maximum likelihood averaged over all possible values of the sufficient statistic (or, more intuitively, if it performs ‘better than average’ on the observed sufficient statistic). Remarkably, this is in open contrast with a comparison based only on the likelihoods, according to which the microcanonical formulation should always be preferred, independently of the observed data.
Expression (III.3) shows that the entity of the difference between the two description lengths crucially depends on the value of the KL divergence. As said in the introduction, the canonical and the microcanonical ensembles are (non-)equivalent if the relative $\text{D}_{\text{KL}}$ is (non-)vanishing as the system size approaches infinity. When studying matrices, identifying the system size with the number of entries, i.e. $nm$, may be inappropriate as adding (only) one entry to a matrix without altering its structure is, in fact, impossible. In these cases, it may be much more reasonable to identify the system size with the number $n$ of items: increasing the system size from $n$ to $n+1$ would, thus, correspond to adding a node to a network or a time-point to a time series.
In this scenario, the canonical and the microcanonical ensembles are equivalent if
$\lim_{n\to\infty}\frac{D_{\text{KL}}(\mathbf{c}^{*})}{n}=\lim_{n\to\infty}\frac{\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\rvert}{n}=0$ (35)
or, equivalently, if
$D_{\text{KL}}(\mathbf{c}^{*})=\lvert\Delta\ln\mathcal{L}(\mathbf{c}^{*})\rvert=o(n).$ (36)
where $o(x)$ stands for a quantity that goes to 00 when divided by $x$ as $n\to+\infty$.
This definition of non-equivalence relies on comparing the likelihood functions of the two models. We aim to extend it in order to take into account the models’ complexities. As a first step, we consider the order of $\Delta\text{DL}$. Notice that
$D_{\text{KL}}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}\quad\Rightarrow\quad\Delta\text{DL}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}$ (37)
as it follows directly from (III.3). Thus, if the two ensembles are equivalent, the difference in description lengths between the corresponding models will be $o(n)$. However, a similar conclusion cannot be immediately derived when non-equivalence holds and $D_{\text{KL}}$ grows at least linearly with $n$. Indeed, the likelihood and complexity terms in (III.3) are of the same order and have opposite signs; therefore, evaluating the order of $\Delta\text{DL}$ is a non trivial task in this case. The following section addresses this problem by explicitly evaluating $\Delta\text{DL}$ in two notable examples of maximum-entropy models.
IV Applications to maximum-entropy models
The simplest constraint one may imagine to enforce is the sum of the elements of a matrix, i.e.
$l(\mathbf{G})\equiv\sum_{i=1}^{n}\sum_{j=1}^{m}g_{ij}.$ (38)
In the case of bipartite binary networks, this quantity corresponds to the total number of links and induces the well-known Erdös-Rényi model. While this model is one of the most popular ones, it is also widely recognized to fall short in capturing the topological heterogeneity that characterizes most real-world systems. An alternative approach to introduce heterogeneity is to constrain the sum of elements of each row, i.e.
$r_{i}(\mathbf{G})\equiv\sum_{j=1}^{m}g_{ij}\quad i=1\dots n.$ (39)
In the case of bipartite binary networks, the sequence $r_{i}(\mathbf{G})$, $i=1\dots n$ coincides with the degree sequence of the nodes belonging to just one layer and induces the canonical MEM known as Bipartite Partial Configuration Model saracco17 . As we will employ asymptotic results to compare the canonical and the microcanonical description lengths, we need to specify the scaling of the constraints $\mathbf{c}^{*}$ as $n\to\infty$. In what follows, we will focus on dense matrices, i.e. we will assume that $g_{ij}^{*}=\Theta(1)$ $\forall(i,j)$, where $\Theta(x)$ indicates a quantity of the same order of $x$ as $n\to\infty$. This choice leads to $r_{i}^{*}=\Theta(m)$ and $l^{*}=\Theta(nm)$, the ‘actual’ order of these quantities depending on the growth rate of $m$. We will consider two cases: (i) $m$ finite in the limit, i.e. $m=\Theta(1)$ and (ii) $m$ growing linearly with $n$, i.e., $m=\Theta(n)$. As a consequence of our assumption, the matrix density
$p^{*}=\frac{l^{*}}{nm}$ (40)
and the row densities
$p_{i}^{*}=\frac{r_{i}^{*}}{m}\quad i=1\dots n$ (41)
will remain finite in the asymptotic limit, i.e. $p^{*}=\Theta(1)$ and $p_{i}^{*}=\Theta(1)$, $\forall\>i$.
IV.1 Enforcing one global constraint
Since we are considering binary matrices, constraining the total number of $1$s in a microcanonical fashion leads to a number of configurations equal to $\Omega(l)=\binom{nm}{l}$. The corresponding microcanonical distribution, thus, becomes
$P_{\text{mic}}(\mathbf{G};l)=\begin{dcases}\frac{1}{\binom{nm}{l}}&\text{if}\>l(\mathbf{G})=l\\ 0&\text{else},\end{dcases}$ (42)
inducing the maximum log-likelihood
$\ln\mathcal{L}_{\textrm{mic}}(l)=-\ln\binom{nm}{l}.$ (43)
The complexity equals the logarithm of the number of graphical values, i.e.
$\text{COMP}_{\text{mic}}^{\text{NML}}=\sum_{l=0}^{nm}1=\ln[nm+1].$ (44)
The microcanonical description length, thus, reads
$\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*})=\ln\binom{nm}{l^{*}}+\ln[nm+1].$ (45)
The canonical probability can be computed explicitly, as this model represents one of the few canonical MEMs whose partition function can be computed analytically. Specifically, it reads $Z(\theta)=(1+e^{\theta})^{nm}$ and induces the expression
$P_{\text{can}}(\mathbf{G};\theta)=\frac{e^{-\theta l(\mathbf{G})}}{(1+e^{-\theta})^{nm}}.$ (46)
Upon defining
$p\equiv\frac{e^{-\theta}}{1+e^{-\theta}},$ (47)
we can re-write the canonical probability via the so-called mean-value parametrization leading to the expression
$P_{\text{can}}(\mathbf{G};p)=p^{l(\mathbf{G})}(1-p)^{nm-l(\mathbf{G})}$ (48)
i.e. the distribution of a collection of $nm$ i.i.d. Bernoulli variables, indicating that each matrix element is either $1$, with probability $p$, or 00, with probability $1-p$. The ML estimator of $p$ is given by the matrix density $\hat{p}(\mathbf{G})=l(\mathbf{G})/nm$ and the maximum log-likelihood, obtained by inserting $\hat{p}$ into (48), reads
$\ln\mathcal{L}_{\textrm{can}}(l)=-nm\cdot h\left(\frac{l}{nm}\right)$ (49)
where $h(p)=-p\ln(p)-(1-p)\ln(1-p)$ is the entropy of a Bernoulli distribution with a parameter equal to $p$. An exact expression for the canonical complexity has been derived in staniczenko and reads
$\text{COMP}_{\textrm{can}}^{\text{NML}}=\ln\left[\frac{e^{nm}\Gamma(nm,nm)}{nm^{nm-1}}+1\right]$ (50)
where $\Gamma(s,x)$ is the upper incomplete gamma function. In case $s$ is a positive integer (as it happens to be our case), it can be expressed as
$\Gamma(s,x)=e^{-x}(s-1)!\sum_{k=0}^{s-1}\frac{x^{k}}{k!}.$ (51)
In conclusion, the canonical description length reads
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\ln\left[\frac{e^{nm}\Gamma(nm,nm)}{nm^{nm-1}}+1\right].$ (52)
Notably, the details of the matrix structure play no role as $\text{DL}_{\text{can}}^{\text{NML}}$ solely depend on global quantities, namely the number of $1$s and the sample size - in our case, $nm$.
So far, all the reported results are exact. Nevertheless, as we are interested in the order of $\Delta\text{DL}$ when $n\to\infty$, we consider the following asymptotic results, derived from the exact ones (see A.1):
$\Delta\ln\mathcal{L}(\mathbf{G^{*}})$ $=-\frac{1}{2}\ln[nm]-\frac{1}{2}\ln[2\pi p^{*}(1-p^{*})]+o(1),$ (53)
$\Delta\text{COMP}$ $=-\frac{1}{2}\ln[nm]+\frac{1}{2}\ln\left[\frac{\pi}{2}\right]+o(1)$ (54)
and, upon combining them, the (asymptotic) difference between description lengths becomes
$\Delta\text{DL}(\mathbf{G}^{*})=\frac{1}{2}\ln[\pi^{2}p^{*}(1-p^{*})]+o(1).$ (55)
The expressions above provide two significant insights. First, the order of the difference between log-likelihood functions is $o(n)$, regardless of the growth rate of $m$, a result implying that condition (36) is met and ensemble equivalence holds. Second, the leading terms of expressions (53) and (54) are identical, thus cancelling out when compared. Consequently, the (asymptotic) difference between description lengths depends on the size of the system only through the matrix density $p^{*}$, assumed to be finite in the dense regime. Thus, in the same regime, such a difference is finite, in agreement with (37).
IV.2 Enforcing $n$ local constraints
Let us now consider canonical and microcanonical models induced by an extensive number of parameters, each one corresponding to the sum of the elements of a different row. With this choice, the matrix can be viewed as a collection of $n$ independent strips of size $1\times m$. Since we impose a single constraint per row, all the prior results hold row-wise, with $m$ replacing $nm$ and $r_{i}$ replacing $l$. Specifically, the microcanonical probability distribution reads
$P_{\text{mic}}(\mathbf{G};\mathbf{r})=\begin{dcases}\prod_{i=1}^{n}\frac{1}{\binom{m}{r_{i}}}&\text{if}\>\mathbf{r}(\mathbf{G})=\mathbf{r}\\ 0&\text{else}\end{dcases}$ (56)
and the related maximum log-likelihood is the sum of the individual ones, i.e.
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{G}^{*})=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}^{*}}.$ (57)
Analogously, the microcanonical complexity reads
$\text{COMP}_{\textrm{mic}}=n\ln[m+1]$ (58)
inducing the expression
$\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*})=\sum_{i=1}^{n}\ln\binom{m}{r_{i}^{*}}+n\ln[m+1].$ (59)
Similarly, the canonical probability is the product of row-specific canonical probabilities
$P_{\text{can}}(\mathbf{G};\bm{\theta})=\prod_{i=1}^{n}\frac{e^{-\theta_{i}r_{i}(\mathbf{G})}}{(1+e^{-\theta_{i}})^{m}}$ (60)
and can be re-written as
$P_{\text{can}}(\mathbf{G};\mathbf{p})=\prod_{i=1}^{n}p_{i}^{r_{i}(\mathbf{G})}(1-p_{i})^{m-r_{i}(\mathbf{G})}$ (61)
upon considering the mean-value parametrization
$p_{i}\equiv\frac{e^{-\theta_{i}}}{(1+e^{-\theta_{i}})^{m}}\quad i=1\dots n.$ (62)
Since for each row the ML estimator of the parameter $p_{i}$ is given by the row density $\hat{p}_{i}(\mathbf{G})=\frac{r_{i}(\mathbf{G})}{m}$, the maximum log-likelihood reads
$\ln\mathcal{L}_{\textrm{can}}(\mathbf{r})=-m\sum_{i=1}^{n}h\left(\frac{r_{i}}{m}\right)$ (63)
and the complexity reads
$\text{COMP}_{\textrm{can}}=n\ln\left[\frac{e^{m}\Gamma(m,m)}{m^{m-1}}+1\right].$ (64)
In conclusion, the canonical description length reads
$\text{DL}_{\text{can}}(\mathbf{G}^{*})=m\sum_{i=1}^{n}h\left(\frac{r_{i}}{m}\right)+n\ln\left[\frac{e^{m}\Gamma(m,m)}{m^{m-1}}+1\right].$ (65)
In the hypothesis of $m$ growing linearly with $n$, i.e. $m=\Theta(n)$, we consider the following asymptotic expressions (see A.2)
$\Delta\ln\mathcal{L}(\mathbf{G^{*}})=$ $-\frac{n}{2}\ln[m]-\frac{1}{2}\sum_{i=1}^{n}\ln[2\pi p_{i}^{*}(1-p_{i}^{*})]$
$+\frac{n}{12m}-\frac{1}{12m}\sum_{i=1}^{n}\frac{1}{p_{i}^{*}(1-p_{i}^{*})}+o(1),$ (66)
$\Delta\text{COMP}=$ $-\frac{n}{2}\ln[m]+\frac{n}{2}\ln\left[\frac{\pi}{2}\right]$
$+a\cdot\frac{n}{\sqrt{m}}+b\cdot\frac{n}{m}+o(1)$ (67)
with
$a$ $=\frac{2}{3}\sqrt{\frac{2}{\pi}}\simeq 0.53,$ (68)
$b$ $=-\frac{11}{12}-\frac{4}{9\pi}\simeq-1.06.$ (69)
By combining these expressions, we obtain the asymptotic expression
$\Delta\text{DL}^{\text{}}(\mathbf{G}^{*})=$ $\frac{1}{2}\sum_{i=1}^{n}\ln[\pi^{2}p_{i}^{*}(1-p_{i}^{*})]+a\cdot\frac{n}{\sqrt{m}}$
$+c\cdot\frac{n}{m}+\frac{1}{12m}\sum_{i=1}^{n}\frac{1}{p_{i}^{*}(1-p_{i}^{*})}+o(1)$ (70)
where $c=-1-\frac{4}{9\pi}\simeq-1.14$.
As for the case of one global constraint, we focus on two aspects of these results. First, the difference between log-likelihood functions is either $\Theta(n)$ or $\Theta(n\ln n)$ depending on the order of $m$, i.e. $m=\Theta(1)$ in the first case and $m=\Theta(n)$ in the second case. In any case, according to (36), the ensemble equivalence is broken. In particular, if $m=\Theta(n)$, the leading terms of expressions (IV.2) and (IV.2) are both of order $\Theta(n\ln n)$, thus cancelling out when compared - precisely as in the equivalent case. Nonetheless, the remaining terms are still of order $\Theta(n)$. Hence, the (asymptotic) difference between description lengths increases linearly with the system size, regardless of the growth rate of $m$.
IV.3 Future research: Towards model-level equivalence
In this section, we explicitly show that, in both cases considered, the difference between likelihoods and the difference between complexities share the exact same leading order. This confirms that the contribution of the complexities in comparing the two formulations is as important as that of the likelihoods.
Furthermore, relation (37) is confirmed in the case of one global constraint examined here, where equivalence holds and the difference between the canonical and microcanonical description lengths turns out to be $o(n)$. At the same time, the result that $\Delta\text{DL}$ is $\Theta(n)$ in the case of $n$ local constraints shows that the signature of the breaking of ensemble equivalence is still visible when the models’ complexities are involved in the comparison.
These findings offer valuable insights for future research on non-equivalence between microcanonical and canonical MEMs within the model selection framework. They lead to the following question: can we generalize the definition of equivalence at the level of the models, i.e. by including their complexities and independently of the observed data?
$\Delta\text{DL}$ is not suitable for this purpose: while it does take into account the models’ complexities, it still depends on the observed values. Instead, we propose a definition involving the universal NML distributions representing the two models. Inspired by definition 35, we consider the following KL divergence
$D^{\text{NML}}_{\text{KL}}$ $\equiv S(P_{\text{mic}}^{\text{NML}}||P_{\text{can}}^{\text{NML}})$ (71)
$=\sum_{\mathbf{G}\in\mathcal{G}}\bar{P}^{\text{NML}}_{\textrm{mic}}(\mathbf{G})\ln\frac{\bar{P}^{\text{NML}}_{\textrm{mic}}(\mathbf{G})}{\bar{P}^{\text{NML}}_{\textrm{can}}(\mathbf{G})}$
$=\sum_{\mathbf{G}\in\mathcal{G}}\bar{P}^{\text{NML}}_{\textrm{mic}}(\mathbf{G})\Delta\text{DL}(\mathbf{G})$
$=\sum_{\mathbf{c}\in\mathcal{C}}\Omega(\mathbf{c})\frac{1}{\Omega(\mathbf{c})\mathcal{N}_{\mathcal{C}}}\Delta\text{DL}(\mathbf{c})$
$=\overline{\Delta\text{DL}(\mathbf{c})},$
where $\overline{\Delta\text{DL}(\mathbf{c})}$ represents the uniform average of $\Delta\text{DL}(\mathbf{c})$ over all possible values of $\mathbf{c}\in\mathcal{C}$. Following the standard definition and assuming that the more relevant scale to consider is still the system size $n$, we say that the canonical and microcanonical models are equivalent on the model level if
$\lim_{n\to\infty}\frac{D^{\text{NML}}_{\text{KL}}}{n}=\lim_{n\to\infty}\frac{\overline{\Delta\text{DL}(\mathbf{c})}}{n}=0.$ (72)
Notice that, according to (37)
$D_{\text{KL}}(\mathbf{c})=o(n)\quad\forall\mathbf{c}\in\mathcal{C}\quad\Rightarrow\quad D_{\text{KL}}^{\text{NML}}=o(n).$ (73)
This relation states that whenever ensemble equivalence holds, the corresponding microcanonical and canonical models are equivalent on the model level. The reverse implication, verified here in the specific case of one-sided local constraints, will be investigated in future work.
V NML-based VS Bayesian approach to model selection
In the previous sections, we implicitly assume that no prior knowledge about the parameters is available. Under this assumption, the NML-based universal distribution is the best choice in a precise minimax sense, as shown in section II. Here, we consider the Bayesian approach to model selection grunwald_book ; grunwald2019 by introducing the Bayesian universal distribution
[…]
(see  peixoto17 ; vallescatala18 for applications in the context of network models). The main difference with respect to the NML description length lies in the presence of a prior on the parameters $w(\bm{\theta})$ to be supplied by the user. This implementation, which is formally equivalent to the Bayes factor method kass95 , provides a description length for $\bm{x}$ through model $\mathcal{M}$ that is defined as
$\text{DL}_{\mathcal{M}}^{\text{Bayes}}(\bm{x})\equiv-\ln\bar{P}^{\text{Bayes}}_{\mathcal{M}}(\bm{x}).$ (75)
In some cases, the NML-based approach can be retrieved in a Bayesian context once the Bayesian distribution is equipped with a prior such that the two description lengths are, at least asymptotically, the same. We will refer to these priors as NML-optimal priors. In the microcanonical case, the uniform prior is an NML-optimal prior. Indeed, when the parameter space is a discrete set, the integral in (75) is replaced by a sum, and the microcanonical Bayesian description length reads
$\text{DL}_{\text{mic}}^{\text{Bayes}}(\mathbf{G}^{*})$ $=-\ln\mathcal{L}_{\text{mic}}(\mathbf{G}^{*})-\ln w(\mathbf{c}^{*})$
$=\ln\Omega(\mathbf{c}^{*})-\ln w(\mathbf{c}^{*}),$ (76)
which coincides with the NML-based description length provided by equation (19), upon identifying $w(\mathbf{c})$ with a uniform prior over $\mathcal{C}$:
$w^{\text{Uniform}}(\mathbf{c})=\frac{1}{\mathcal{N}_{\mathcal{C}}}\quad\forall\>\mathbf{c}\in\mathcal{C}.$ (77)
In the canonical case, the situation is quite different. First, consider $V$ i.i.d. outcomes $\bm{x}=(x_{1},\dots x_{V})$ from a general $k$-dimensional exponential model. When $\bm{x}$ is such that the ML estimators of $\bm{\theta}$ are bounded away from the boundaries of $\Theta$ as the sample size goes to infinity, and under some regularity conditions on $\mathcal{M}$ which hold in our setting, the following asymptotic results hold true:
$\text{DL}^{\textrm{NML}}_{\mathcal{M}}(\bm{x})=$ $-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}$
[…]
$\text{DL}_{\mathcal{M}}^{\text{Bayes}}(\bm{x})=$ $-\ln\mathcal{L}_{\mathcal{M}}(\bm{x})+\frac{k}{2}\ln\frac{V}{2\pi}$
$-\ln w(\hat{\bm{\theta}}(\bm{x}))+\ln\sqrt{\text{det}\mathbf{I}(\hat{\bm{\theta}}(\bm{x}))}+o(1),$ (79)
where $\text{det}\mathbf{I}(\theta)$ is the determinant of the $k\times k$ normalized Fisher information matrix. If we equip the Bayesian universal distribution with the Jeffreys prior jeffreys1946
[…]
by posing $w(\bm{\theta})\equiv J(\bm{\theta})$, the two asymptotic expressions coincide (for a formal derivation of these results, we redirect the interested reader to  rissanen96 and  grunwald_book ). Thus, whenever the asymptotic formulas (V) and (V) hold true, the NML-based description length is asymptotically equivalent to the one provided by the Bayesian distribution equipped with Jeffreys prior (hereby, Bayes-Jeffreys description), which is NML-optimal.
This is verified in the case of one global constraint, where the resulting canonical model turns out to be an exponential i.i.d. model. The asymptotic formulas do not provide the rate of convergence between the two description lengths, which can be derived by directly comparing the exact expressions of the NML description length $\text{DL}_{\text{can}}^{\text{NML}}$ (65) and the Bayes-Jeffreys description length
$\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\ln[\pi(nm)!]\\ -\ln\left[\Gamma(nm-l^{*}+1/2)\Gamma(l^{*}+1/2)\right]$ (81)
(see B.1). In the limit $n\to\infty$ one finds
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+O\left(\frac{1}{nm}\right).$ (82)
The same comparison can be carried out in the case of $n$ local constraints by applying the results above to each row. In this case, the Bayes-Jeffreys description length reads
$\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=n\ln[\pi m!]\\ -\sum_{i=1}^{n}\ln\left[\Gamma(m-r_{i}^{*}+1/2)\Gamma(r_{i}^{*}+1/2)\right]$ (83)
and in the limit $n\to\infty$ one finds
$\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*})-\text{DL}_{\text{can}}^{\text{Bayes-Jeffreys}}(\mathbf{G}^{*})=\frac{2}{3}\frac{\sqrt{2}n}{\sqrt{\pi m}}+\Theta\left(\frac{n}{m}\right).$ (84)
As $n\to\infty$, the last difference diverges both in case $m$ stays finite, i.e. $m=\Theta(1)$, and in case $m$ diverges, upon assuming that the rate at which $m$ diverges does not exceed that of $n$. Consequently, in this case, it is no longer true that the NML approach can be asymptotically retrieved from the Bayesian one equipped with Jeffreys prior. Although the existence of a modified Jeffreys prior extending the equivalence of the two approaches to the case of $n$ local constraints can indeed be hypothesized, for the time being, we merely suggest not to assume this to be true when ensemble non-equivalence holds.
In the remainder of this section, we show that the Bayesian description length becomes more sensitive to the prior choice whenever local constraints are enforced. Let us start by considering the relationship
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}$ (85)
stating that the NML-based description length of a canonical model coincides with the Bayesian description length of its microcanonical variant, equipped with the so-called canonical prior, introduced by Rissanen in rissanen2001 and defined in B.2. Moreover,
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}}$ (86)
i.e. the NML-based description length of the microcanonical model coincides with the Bayesian description length of its canonical variant, equipped with a uniform prior on the mean-value parameters.
The two identities, derived in B.2, show that there exists a prior guaranteeing that the Bayesian description length of the canonical (microcanonical) model coincides with the NML-based description length of the microcanonical (canonical) model. This might lead to the conclusion that, in the framework of model selection, canonical and microcanonical models coincide even when non-equivalence holds if the right priors are chosen. Nevertheless, these priors are not NML-optimal. While this non-optimality is negligible in the case of one global constraint, it could play a crucial role when local constraints are enforced. In other words, in this case, different choices of prior can result in very different Bayesian description lengths for microcanonical as for canonical models.
The $n$ local constraints case examined in the previous sections provides an instructive example. In this case, we showed that the difference between the canonical and microcanonical NML description lengths is $\Theta(n)$. Combining this result with (85), we can conclude that when it comes to the microcanonical model, choosing between the uniform prior (which is NML-optimal) and the canonical prior can amount to a difference in description length of $\Theta(n)$ nats. Similarly, when it comes to the canonical model, it can be easily shown that, based on (84, 86) when $m=\Theta(n)$, choosing between the uniform and the Jeffreys prior can lead to a difference in description length of $\Theta(\sqrt{n})$ nats. In contrast, the same differences are $\Theta(1)$ when one global constraint is involved.
This simple example shows that when non-equivalence holds, the selection of different priors can result in increasingly larger differences in DL when the system size grows.
VI Conclusion
This study examined the implications of ensemble non-equivalence within MDL model selection. Specifically, we compared the NML-based description lengths of canonical and microcanonical maximum-entropy models.
We found that while microcanonical models fit the data better than their canonical counterparts, they are also more complex. The model achieving the minimum description length depends on the observed values of the sufficient statistics.
We explicitly compared the description lengths of canonical and microcanonical models of dense $n\times m$ binary matrices in a couple of cases of interest: one global constraint, such as the sum of the entries of the matrix, and $n$ local constraints, such as the sum of the entries of each row of the matrix. While the former induces a homogeneous model, according to which the probability of an entry being 1 is the same for all entries, the latter induces a heterogeneous model, according to which the probability of an entry being 1 depends on the specific row. In the second scenario, microcanonical and canonical ensembles are no longer equivalent. By comparing the description length of the canonical and the microcanonical formulations of these two models, we observed that their difference remains finite when one global constraint is enforced while growing linearly with the system size when $n$ local constraints are enforced.
Our findings shed light on the consequences that may arise from choosing either hard or soft constraints in terms of data compression, leading us to conclude that canonical and microcanonical models should be regarded as distinct from a model selection perspective in the presence of ensemble non-equivalence. Moreover, they suggest a possible new definition of model equivalence based on MDL that needs further investigation.
Furthermore, we explored the relationship between NML-based and Bayesian description lengths. While the Jeffreys prior is NML-optimal for canonical models under a single global constraint, this is no longer true when local constraints are enforced. In this case, the difference between these two description lengths diverges as the system size increases. This observation should warn practitioners regarding the two approaches as interchangeable in the presence of ensemble non-equivalence, and it raises the question of whether an NML-optimal prior exists when local constraints are enforced.
Lastly, we showed that different priors can result in largely different description lengths when local constraints are involved. This outcome underscores the importance of carefully selecting priors when dealing with an extensive number of constraints.
Appendix A Asymptotic expansions
Here, we derive the asymptotic expansions used to obtain the asymptotic results in section IV.
A.1 One global constraint
We start by deriving an asymptotic expansion for the microcanonical log-likelihood:
$\ln\mathcal{L}_{\text{mic}}(l)=-\ln\binom{nm}{l}=-\ln(nm)!+\ln(nm-l)!+\ln l!.$ (87)
In the dense regime, $l$ is of order $\Theta(n)$ and Stirling’s approximation
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\Theta(1/x)$ (88)
can be used to evaluate all the factorials above, yielding
$\ln\mathcal{L}_{\text{mic}}(l)=-nm\ln(nm)+(nm-l)\ln(nm-l)+l\ln l-\frac{1}{2}\ln\left(\frac{nm}{2\pi l(nm-l)}\right)+o(1),$ (89)
which can be further expressed as a function of the density $p=l/nm$:
$\ln\mathcal{L}_{\text{mic}}(p)=-nm\cdot h(p)+\frac{1}{2}\ln(nm)+\frac{1}{2}\ln(2\pi p(1-p))+o(1).$ (90)
By combining this expression with the canonical log-likelihood
$\ln\mathcal{L}_{\text{can}}(p)=-nm\cdot h(p),$ (91)
we obtain the asymptotic log-likelihood difference of equation (53).
Similarly, we combine the asymptotic expansion of the microcanonical complexity
$\text{COMP}_{\text{mic}}$ $=\ln(nm+1)=\ln(nm)+\ln\left(1+\frac{1}{nm}\right)$ (92)
$=\ln(nm)+o(1),$
which follows from $\ln\left(1+\frac{1}{x}\right)=\Theta(1/x)$, and the following asymptotic expansion of the canonical complexity
$\text{COMP}_{\text{can}}=\frac{1}{2}\ln(nm)+\frac{1}{2}\ln\left(\frac{\pi}{2}\right)+o(1).$ (93)
The latter is easily derived by asymptotically expanding the following approximation
$\text{COMP}_{\textrm{can}}^{\text{NML}}\simeq\ln\left[\sqrt{\frac{\pi nm}{2}}+\frac{2}{3}+\frac{\sqrt{2\pi}}{24\sqrt{nm}}-\frac{4}{135nm}+\frac{\sqrt{2\pi}}{576\sqrt{(nm)^{3}}}+\frac{8}{2835(nm)^{2}}\right],$ (94)
provided in  [25], where the authors show it to be very precise already for small values of the total number of entries $nm$. Finally, equation (54) is obtained by comparing the two asymptotic complexities. Notice that the asymptotic formula (93) for the canonical complexity represents a very well-known result, as it can be derived directly by applying the asymptotic formula (V) to the case of i.i.d. Bernoulli random variables.
A.2 One-sided local constraints
Here, we consider the regime in which $m=\Theta(n)$. Similarly to the previous case, we need to expand the microcanonical likelihood
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-\sum_{i=1}^{n}\ln\binom{m}{r_{i}}=-n\ln m!+\sum_{i=1}^{n}\ln(m-r_{i})!+\ln r_{i}!.$ (95)
In the dense regime, all $r_{i}$’s are of order $\Theta(n)$, and we could again apply Stirling’s formula to the factorials above. Nevertheless, because of the factor $n$ ahead of everything, Stirling’s formula is not enough, and we turn to Stirling’s series [31]
$\ln x!=x\ln x-x+\frac{1}{2}\ln(2\pi x)+\frac{1}{12x}+\Theta(1/x^{3}).$ (96)
The resulting asymptotic expression can be expressed as a function of the row densities $p_{i}=r_{i}/m$:
$\ln\mathcal{L}_{\textrm{mic}}(\mathbf{r})=-m\sum_{i=1}^{n}h(p_{i})+\frac{n}{2}\ln m+\frac{1}{2}\sum_{i=1}^{n}[\ln(2\pi p_{i}(1-p_{i}))]\\ -\frac{n}{12m}+\frac{1}{12m}\sum_{i=1}^{n}\left[\frac{1}{p_{i}(1-p_{i})}\right]+o(1).$ (97)
By combining this expression with the canonical log-likelihood
$\ln\mathcal{L}_{\text{can}}(\mathbf{p})=-m\sum_{i=1}^{n}h(p_{i}),$ (98)
one obtains the asymptotic log-likelihood difference of equation (IV.2).
Similarly, the asymptotic expansion of the microcanonical complexity
$\text{COMP}_{\text{mic}}$ $=n\ln(m+1)=\ln m+n\ln\left(1+\frac{1}{m}\right)$
$=n\ln m+\frac{n}{m}+o(1),$ (99)
which follows from $\ln\left(1+\frac{1}{x}\right)=\frac{1}{x}+\Theta(1/x^{2})$, is compared to the following asymptotic expansion of the canonical complexity, based on (94)
$\text{COMP}_{\text{can}}=\frac{n}{2}\ln m+\frac{n}{2}\ln\frac{\pi}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi}}\cdot\frac{n}{\sqrt{m}}+\left(\frac{1}{12}-\frac{4}{9\pi}\right)\cdot\frac{n}{m}+o(1)$ (100)
to express (64). The asymptotic expression (IV.2) is obtained as the difference between the two asymptotic complexities.
Appendix B NML and Bayes
B.1 Bayesian-Jeffreys description lengths
In what follows, we compute the Bayesian description length $\text{DL}^{\text{B, Jeffreys}}_{\text{can}}$ of the canonical model obtained by constraining the sum $l$ over the matrix. As already stated, this model is equivalent to modelling $nm$ i.i.d. Bernoulli variable, and this result can be found in Example 8.3 of [18]. Later, we extend this result to the case of $n$ one-sided local constraints.
First, we compute Jeffreys prior according to (80). For a Bernoulli variable, the Fisher information for the parameter $p$ is $\mathbf{I}(p)=p^{-1}(1-p)^{-1}$ and $\int_{0}^{1}\sqrt{\mathbf{I}(p)}=\pi$. Thus, the Jeffreys prior reads
$J(p)=\frac{1}{\pi\sqrt{p(1-p)}}$ (101)
and, according to definition (75), the Bayesian-Jeffreys DL is
[…]
$=-\ln\int_{0}^{1}p^{l^{*}}(1-p)^{nm-l^{*}}\frac{1}{\pi\sqrt{p(1-p)}}$
[…]
The integral above is the Beta function
[…]
computed for $x=l^{*}+\frac{1}{2}$ and $y=nm-l^{*}+\frac{1}{2}$, where $\Gamma(x)$ is the Gamma function
[…]
Thus
$\text{DL}^{\text{B, Jeff}}_{\text{can}}(\mathbf{G}^{*})=-\ln\frac{\Gamma\left(l^{*}+\frac{1}{2}\right)\Gamma\left(nm-l^{*}+\frac{1}{2}\right)}{\pi(nm)!},$ (105)
which corresponds to Equation (81). By employing the Stirling series, this expression can be expanded asymptotically as
$\text{DL}^{\text{B, Jeff}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+O\left(1/nm\right).$ (106)
The last expression is compared with the asymptotic expansion of the canonical NML description length $\text{DL}^{\text{NML}}_{\text{can}}$, obtained as the difference between the asymptotic expansion of (94) and the canonical log-likelihood:
$\text{DL}^{\text{NML}}_{\text{can}}(\mathbf{G}^{*})=nm\cdot h\left(\frac{l^{*}}{nm}\right)+\frac{1}{2}\ln\frac{\pi nm}{2}+\frac{2}{3}\sqrt{\frac{2}{\pi nm}}+o(1).$ (107)
Finally, Equation (82) is obtained as the difference between (107) and (106)
B.2 Proving identities
In what follows, we prove identities (85) and (86) of section V.
Proving Identity (85).
We begin with identity (85), namely:
$\text{DL}_{\text{can}}^{\text{NML}}=\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}},$ (108)
where the canonical prior $\hat{W}$ inducing the Bayes-Rissanen description length is expressed as
$\hat{W}(\mathbf{c})$ $=\frac{\sum_{\mathbf{G}:\mathbf{c}(\mathbf{G})=\mathbf{c}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G})}{\sum_{\mathbf{G}}P_{\text{can}}(\mathbf{G};\hat{\bm{\theta}}(\mathbf{G}))}$ (109)
$=\frac{\Omega(\mathbf{c})\mathcal{L}_{\text{can}}(\mathbf{c})}{\sum_{\mathbf{G}}\mathcal{L}_{\text{can}}(\mathbf{G})}.$
We recognize the canonical complexity in the denominator of $\hat{W}$. Thus, putting this expression in the Bayesian microcanonical description length, we get
$\text{DL}_{\text{mic}}^{\text{Bayes-Rissanen}}(\mathbf{G}^{*})$ $=\ln\Omega(\mathbf{c}^{*})-\ln\hat{W}(\mathbf{c}^{*})=$ (110)
$=-\ln\mathcal{L}_{\text{can}}(\mathbf{c})+\text{COMP}_{\text{can}}$
$=\text{DL}_{\text{can}}^{\text{NML}}(\mathbf{G}^{*}),$
which proves identity (85).
Proving Identity (86).
We will now prove identity (86), namely:
$\text{DL}_{\text{mic}}^{\text{NML}}=\text{DL}_{\text{can}}^{\text{Bayes-Uniform}},$ (111)
for the specific MEMs considered in this work, starting from the case of one global constraint. The uniform prior on the mean-value parameter $p$ reads
$w^{\text{Uniform}}(p)=1\quad\text{for }p\in[0,1].$ (112)
Putting this prior in the Bayesian description length yields:
[…]
$=\ln\binom{nm}{l^{*}}+\ln(nm+1)$
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$
where the integral is computed by integrating by parts $l^{*}$ times. This proves identity (86) for the case of one global constraint.
The identity holds as well for the case of $n$ one-sided local constraint, with the uniform prior reading
$w^{\text{Uniform}}(\mathbf{p})=1\quad\text{for }\mathbf{p}\in[0,1]^{n}$ (114)
Indeed, we have that
[…]
$=\sum_{i=1}^{n}\ln\binom{m}{r_{i}^{*}}+n\ln(m+1)$
$=\text{DL}_{\text{mic}}^{\text{NML}}(\mathbf{G}^{*}),$ (116)
which proves identity (86).